# Class-Discriminative Attention Maps for Vision Transformers

### v1

@@ 4 Results
Figure 3: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2). The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept. In each instance, from left to right: Sample image, AM, and CDAM. Orange and blue correspond to positive and negative values, respectively.
@@
For our experiments, we have been using the ViT-S/8 architecture trained with DINO as a backbone [1], which is available as a pre-trained model in their repository [^2] https://github.com/facebookresearch/dino.
4.1 Relevance Scores from Classifier
@@
To demonstrate our method, we have trained a linear classifier on the ImageNet [33] that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2. Notice how they clearly distinguish the targeted class with positive values (evidence) from the rest of the image, which is assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in AM. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class. This is a nice feature since CDAM inherits the high-quality object segmentations present in AM, which are focused on the objects of interest. We discuss this operating characteristic in Section 5.
One of the most interesting features of CDAM is that it is highly sensitive to the chosen class (Fig. 4, left). For example, when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values. Interestingly, the green stems of other vegetables are also assigned positive values. CDAM reveals that the ViT has learned to associate input features with chosen classes that are largely consistent with human expectations. Additional examples of CDAM and other interpretability methods are showned in the Appendix A (Figs. 7, 8, 9 and 10).
Figure 4: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects. Note that green parts of vegetables are shown to activate the zuccini class (left). The wrinkles in the elephant’s trunk appear to be related to the zebra concept, probably semantically mistaken by the model as zebra stripes (zoom in for a better view).
4.2 Relevance Scores from Similarity Measure
@@
Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest. These images were randomly selected from the corresponding classes in the ImageNet validation set [33] and the CDAM in Fig. 3 were obtained using Eq. 4. Again we observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image. For instance, for the concept hammer, CDAM highlights not only the hammer itself but also the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept. Unrelated tools or parts appear in blue indicating that the model deems them to be not similar to the concept hammer.
@@
Just as CDAM is sensitive to the chosen class, it also produces attention maps that are specific to the chosen concept (Fig. 4, right). This makes the proposed method particulary valuable for understanding the representations the ViT has learned since the explanations for different concepts can give a complementary view. For instance, CDAM shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig. 4, right) which is at first counter-intuitive. However, when compared with the concept zebra, we can see that the model deems this part of the trunk to be similar to a zebra and it probably mistakes the trunk’s wrinkles for the zebra’s stripes. By choosing a set of images with a common concept, CDAM helps probing the latent space of the ViT, allowing us to learn about how it “sees” the world in a semi-supervised manner.
4.3 Comparison with relevance propagation and token ablation maps
Figure 5: Comparison of specificity to the target class. AM is, by definition, not influenced by it, but relevance propagation [2] does show some specificity. In particular due to the signed relevance scores both CDAM and TAM exhibit strong class-sensitivity and the two maps roughly agree, with CDAM appearing as stronger regularized.
Upon visual inspection, it appears that CDAM is estimating patch token relevance that aligns well with our expectation concerning the most relevant (counter-)evidence for a given target class (Figs. 1 and 4). To investigate whether these relevance scores correspond to the actual inference performed by the model, we compare them to token ablation maps (TAMs) and relevance propagation (RP) [2]. In TAMs, the model output is measured after an individual token is removed (Section 3.3). RP, which is based on the Deep Taylor Decomposition, is an explanation method for ViTs that, to the best of our knowledge, produces relevance scores that are the most class-discriminative so far [2, 23].
We find that RP is more class-discriminative than AM and that CDAM and TAM distinguish the different classes more clearly than RP ( Fig. 5). Visually, the signed scores in CDAM and TAM make stronger distinctions between the different classes, whereas RP only “dims” the object that is not targeted. Fig. 6 shows pair-wise comparisons among scores for CDAM, TAM, AM, and RP. We observe that RP tends to assign low but positive scores to tokens that obtained negative scores from CDAM or TAM.
@@ 4.3 Comparison with relevance propagation and token ablation maps
Compared to AM, TAM, and RP, CDAM appears to be less noisy (Fig. 5). We can observe this implicit regularization effect by looking at the scores from 1000 randomly selected images Fig. 6. In AM, TAM, and RP, the per-image average percentage of patch tokens with relevance scores smaller than $5\%$ of the maximum score are $54\%,66\%,49\%$, respectively, whereas in CDAM it is $84\%$. While sparsity may not always imply the most accurate estimation at a single data point (e.g., a pixel), regularization has long been pursued and motivated by interpretability in machine learning [34].
Figure 6: Pairwise scatter plots of relevance/attention scores for CDAM, TAM, AM, and relevance propagation obtained from 1000 randomly selected images from the ImageNet validation set. The target class for the explanations is the one with the highest logit score. Notice in particular that AM and relevance propagation are highly correlated and that CDAM and TAM largely agree on the sign of relevance scores. We also observe that $A_{i}=0\Rightarrow S_{i,c}=0$, where $A_{i}$ and $S_{i,c}$ refer to the AM or CDAM score of token $T_{i}$.
@@ 5 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps (AM), that provides strong discrimination with respect to a chosen class or concept. In contrast, many post-hoc explanation methods produce similar heatmaps even when targeting different classes [23, 17]. CDAM retains appealing characteristics of AM, namely high-quality semantic segmentation in self-supervised ViTs [1]. Previously, a considerable effort has been made to de-noise post-hoc interpretability methods by averaging variations of multiple saliency maps [13, 15]. Without further processing or averaging, the proposed CDAM produces comparatively sparse heat maps that are less noisy than previous methods.
Besides explaining predictions of the classifier on the top of ViT, we also introduce how to obtain a CDAM by measuring the similarity between the latent representation of a sample image and a concept embedding. By probing the latent space of the ViT with different concepts (Fig. 4) we can learn a lot about the representations it has learned. To some degree this can of course also be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [1]. This approach is, however, limited to the categories that the attention heads happened to have learned. CDAM allows us to probe for concepts that are relevant to any given use case.
5.1 Proportionality of CDAM to the attention
CDAM appears to scale the attention $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ by the relevance of the corresponding token $T_{i}$ for a concept. To better understand this observation, without loss of generality, consider a single attention head. Eq. 4 is expanded
@@ 5.1 Proportionality of CDAM to the attention
Importantly, we observe in our multi-head attention ViT model (Fig. 6) that
$A_{i}=0\Rightarrow S_{i,c}=0$ (7)
which implies that in Section 5.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
@@
Empirically, $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$. In other words, CDAM scales $A_{i}$ by the rate at which the dot product (similarity) between $h$ and $c$ changes in the direction of $V_{i}$. In the case that $h$ is just the identity function, we obtain an even simpler interpretation of CDAM. That’s because we then have $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$ and get from Section 5.1, again assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, that $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
5.2 Faithfulness of Post-hoc Explanation Methods
Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT. Therefore, post-hoc explanation methods such as RP, CDAM, and TAM can only explain some facets of the model and can never be comprehensively faithful to it, as that would entail being equivalent to the model itself [17]. Of course, the complete set of the model weights in DNNs would not be considered interpretable. Therefore, attempting to create explanation methods necessitates a balance among many properties, ranging from sparsity to discriminativeness [27].
The proposed CDAM relies on gradients and is therefore a first-order approximation. Instead of referring directly to the input patch tokens, the CDAM scores refer to tokens that, in general, have been enriched with contributions from all other tokens through the attention layers. Relevance is therefore not directly attributed to input features. Nonetheless, by extending the inherent attention scores in ViTs, CDAM captures the high-quality semantic segmentations of input images shown in self-supervised ViTs [1], while providing class-discrimination.
6 Conclusion
By extending the attention maps (AM) inherent to ViT, the proposed CDAM provides relevance scores for a chosen class or concept. The proposed relevance estimation for CDAM essentially scales the attention scores by how relevant they are for a target class or concept. As a result, zero attention scores remain zero in CDAM. The resulting heat maps clearly distinguish the object representing the target from the background, as well as other objects that are present. In addition to using a classifier, we proposed to explain a user-defined concept by selecting a small set of example images. Using similarity measures in the latent representation, CDAM provides highly specific heat maps for a concept that underly selected examples. Especially in the context of SSL where class labels are absent, this makes it a valuable technique to investigate the model representations.
In terms of class-discrimination, our method compares favorably against the state of the art, namely relevance propagation (RP) for transformers [2]. Especially, the positive and negative signs of relevance scores in CDAM make clear distinctions qualitatively and quantitatively. We also devise the token ablation maps (TAM) which visualize the changes in the ViT model output when corresponding image patch tokens are removed. The advantage of TAM is its explicable computation, which serves as a sanity check. CDAM captures the overall patterns of TAM while providing sparser and more semantically consistent heat maps. However, while RP and TAM are used in comparison, neither of them is the “ground truth” of relevance scores.
Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties. Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness, as described in [35]. Note that CDAM is not explicitly designed to satisfy these criteria. For example, we could impose further sparsity constraints that would increase compactness. However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [36]. While this paper uses a ViT model trained with SSL [1], CDAM is applicable more broadly to all transformer models with self-attention, e.g., large language models [37].
Acknowledgments
This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071]. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
Appendix A Appendix
A.1 Multiple attention heads
@@
We note that introducing $h$ attention heads does not change the derivations in Section 5.1 qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$ there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section 5.1 reads
@@
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section 5.1 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
We can also show that CDAM scales the attention by the relevance of the correpsonding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT. Analogously to the above single-head discussion we have
@@
A.2 Examples for explanation methods
Figure 7: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 8: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 9: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 10: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.

### v2

@@ 4 Results
For our experiments using the ImageNet [19], we have been using the ViT-S/8 architecture trained with DINO as a backbone [9], which is available as a pre-trained model in their repository [^1] https://github.com/facebookresearch/dino. We first discuss the results of applying CDAM qualitatively and then perform a quantitative comparison with several baseline methods.
4.1 Qualitative Evaluation
4.1.1 CDAM for a known class
@@
To demonstrate our method, we have trained a linear classifier on the ImageNet [19] that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size of 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2.
CDAM clearly distinguishes the region of interest (ROI) semantically related to the targeted class with positive values (yellow) from the rest of the image. Most backgrounds and other objects are assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in attention maps, see also the scatter plot between CDAM and attention map Fig. 11. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class (Section A.1). This is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps. This qualitative evaluation of shrinkage and sparsity of CDAM is reflected in the quantitative measures for compactness (Section 4.2.1).
One of the most useful features of CDAM is that it is highly sensitive to the chosen class. In Fig. 4 (left), when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values (indicated in blue). Many other importance estimators do not distinguish the target class or only to a lesser degree. In Fig. 6, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad. Relevance propagation demonstrates some selective focus on the ROI, but is not as discriminative as CDAM (Fig. 6). This visual impression of class-sensitivity is supported by the quantitative evaluation in Section 4.2.2. Additional examples of CDAM and other interpretability methods are shown in the Appendix A.
Smooth CDAM appears to reduce the regularization of importance scores, highlighting, for instance, a larger portion of the truck and cars in the background in Fig. 5. This is quite different from how SmoothGrad behaves for CNNs [15], where it usually denoises the importance scores leading to a stronger focus on the ROI. On the other hand, Integrated CDAM can’t, for most samples, be visually distinguished from vanilla CADM even though its performance in the benchmarks is quite different (below).
4.1.2 CDAM for a latent concept
@@
Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest. These 30 images were randomly selected from the corresponding classes in the ImageNet validation set [19] and the CDAM in Fig. 3 were obtained using Eq. 4. We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image. For example, for the concept hammer (bottom right), CDAM primarily highlights the hammer. CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept. Unrelated tools or parts appear in blue indicating that the model deems them to be unrelated to the concept hammer.
@@
Just as CDAM is sensitive to the chosen class, it also produces attention maps that are sensitive to the chosen concept (Fig. 4, right). This makes the proposed method particularly valuable for understanding the representations the ViT has learned, since the explanations for different concepts can give a complementary view. For instance, CDAM distinguishes african elephant and zebra. However, CDAM also shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig. 4, right) which is initially counter-intuitive. However, when targeting the concept zebra, the model deems this part of the trunk to be similar to a zebra. Speculatively, the trunk may be too similar to the zebra’s muzzle or the trunk’s wrinkles may be mistaken for the zebra’s stripes.
4.2 Quantitative evaluation
We have performed quantitative evaluations of the correctness and compactness of our proposed methods, in comparison to multiple well-known importance estimators. We further investigated class sensitivity by using importance scores from multiple classes.
4.2.1 Evaluation of correctness
@@
[…]
4.2.2 Evaluation of class discrimination
[…]
The importance scores for CDAM become anti-correlated with the model output when the wrong importance scores are used for the perturbation (Fig. 10). This anti-correlation probably results from the fact that pixels that are evidence for the correct target class are counter-evidence for the wrong class. This result corroborates the visual impression that CDAM correctly assigns importance scores with opposite signs to the objects corresponding to the targeted class and other objects that are present (Fig. 6). This is also reflected in the negative values of $f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}$ for CDAM (Fig. 9), indicating that the ranking from least to most important has been, at least partially, reversed.
4.2.3 Evaluation of compactness
Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in explainability of deep learning [44]. While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning [47]. Ideally, compact explanations (e.g., sparse and shrunken importance estimators) are preferred, when other properties such as correctness and class sensitivity are constant.
CDAM and Integrated CDAM resulted in the highest degree of compactness, together with Input $\times$ Grad. For those three importance estimators, on average, $88\%$ of importance scores were less than $5\%$ of the maximum score Table 2. IntGrad and SmoothGrad show relatively high sparsity $0.84$ and $80\%$, respectively. But both are lacking in correctness (Section 4.2.1) and class-discrimination (Section 4.2.2). Relevance Propagation shows the lowest sparsity of all considered methods $48\%$, but a high level of correctness. These results suggest that these correctness and compactness metrics are orthogonal.
5 Application to Medical Images
To demonstrate CDAM for another use case we apply it to lung nodule malignancy and biomarker prediction using the LIDC dataset [23].
5.1 Malignancy Prediction
After preprocessing, the LIDC data consists of 443 benign and 411 malignant lung CT scans. Training, validation, and test sets (in the ratios of 0.7225, 0.1275, 0.15) were stratified by and balanced according to these labels, e.g., benign and malignant. We fine-tuned a ViT model, with a DINO backbone pretrained on the ImageNet, for 50 epochs. In a parameter sweep, we varied the number of trainable layers ($10-50$) and dropout rates ($0.0-0.09$), where the learning rate was exponentially decaying ($\alpha=0.0003$ and $\beta=0.95$). The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003222It might be possible to increase the model performance via further investigation. Here, we use a straightforward model and training scheme to demonstrate the application of CDAM on medical images.. Attention maps and CDAM are obtained based on this model.
The obtained attention maps suggest that the model focuses on patches with nodule fragments (Fig. 12(b)). However, we note that attention maps almost always focus on pulmonary nodules without taking into account the downstream classification into benign and malignant samples. CDAM provides detailed structures where positive and negative values are indicated by orange and blue, respectively. In this binary classification where 0 is for benign and 1 for malignancy, input pixels with positive values in CDAMs are driving the classification towards malignancy.
5.2 Biomarker prediction
We turn our attention to clinical biomarkers in LIDC that are routinely used by medical practitioners. Inspired by the concept bottleneck model (CBM) [48], we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter). Essentially, a regression model was built on a pre-trained ViT model with the mean squared error (MSE) as loss function. For investigation of LIDC and incorporation of CBM, see [49].
CDAMs were obtained for 8 biomarkers (Fig. 13). In the context of the margin, patches including the edge of a nodule get positive CDAM scores, whereas patches corresponding to the nodule body get negative CDAM scores. For the diameter, patches corresponding to the nodule body get a positive score, whereas the background pixels exhibit neutral (near zero) or slightly negative scores. When it comes to sphericity, patches including fragments of nodule curvature get a positive CDAM score. Spiculation refers to the presence of needle and spike-like structures, on the margin. High spiculation often is an indication of malignancy. In a majority of cases, CDAM highlights spicules with positive values, whereas the inner bodies of nodules (i.e., not spicules) tend to get low importance scores.
@@ 6 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps, that provides strong discrimination with respect to a chosen class or concept, while exhibiting excellent fidelity and compactness. We introduce additional variants of CDAM, inspired by SmoothGrad ([15]) and IntGrad [16]), which improve the performance on the quantitative benchmarks. Our proposed methods retain appealing characteristics of attention maps, namely high-quality semantic segmentation in self-supervised ViTs. The proposed importance estimation for CDAM scales the attention scores by how relevant they are for a target class or concept (Section A.1). Therefore, zero attention scores remain zero in CDAM and it produces compact explanations with high sparsity and shrinkage.
Post-hoc explanation methods for DNNs aim to make the full decision-making process of the model more transparent by providing an approximation with certain desirable properties, such as correctness, sensitivity, and compactness. Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output [27, 38]. Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps [38]). We find, both qualitatively and quantitatively, that CDAM is highly sensitive to the targeted class, assigning positive importance scores to objects corresponding to the target class and negative ones to other (semantically distinct) objects in the image.
Besides explaining predictions of the classifier on the top of ViT, we introduce several variants. Firstly, by probing the latent space of the ViT (Fig. 4), self-supervised CDAM provides importance scores specific to the user-defined concept. In the context of SSL, where class labels are absent, this makes it a valuable technique to investigate the representations learned by the model, as shared by user-selected samples. To some degree, this can alternatively be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [9]. This approach is, however, limited to the categories that the attention heads happened to have learned. Secondly, we introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs. Notably, instead of adding noise to or modifying the input images as in SmoothGrad [15]) or IntGrad [16], respectively, our methods act on tokens in the final transformer block.
Our application on medical images examplifies the need for fine-grained class-sensitive explanations offered by CDAM. Other explainability methods including attention maps are often highlighting whole nodules and tumors. But they are uninformative about the innerworkings of the ViT model predicting malignancy or biomarkers. While medical implication of these explanations are beyond the scope of this paper, the proof of concept application demonstrates usability of CDAM in cancer images. As the methods of explainable AI (XAI) becomes incorporated into practices, we foresee a greater level of human-computer interaction where class sensitivity will play a critical role for utility of explanations.
Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT. Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties. Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness. Note that CDAM is not explicitly designed to satisfy these criteria. For example, we could impose further sparsity constraints that would increase compactness. However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [50].. While this paper focuses on a ViT model, CDAM is applicable more broadly to all transformer models with self-attention, including large language models (LLMs) [51].
Acknowledgments
This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long-term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071]. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
Appendix A Appendix
A.1 Proportionality of CDAM to attention maps
The proposed class-discriminative attention scores scale the attention scores. Empirically, in all of our evaluation tasks and applications (e.g., Fig. 11), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
$A_{i}=0\Rightarrow S_{i,c}=0.$ (9)
This operating characteristic is highly favorable, as besides introducing class discrimination, CDAM provides implicit regularization resulting in sparser heat maps. Based on a combination of empirical observations and mathematical derivations, we argue that the attention score $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ is scaled by the relevance of the corresponding token $T_{i}$ for a concept or a class.
To better understand this observation, we first consider a concept-based CDAM with a single attention head. Eq. 4 is expanded:
@@ A.1 Proportionality of CDAM to attention maps
Eq. 9 implies that in Section A.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
@@
We thus find that $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$. In other words, CDAM scales $A_{i}$ by the rate at which the dot product (similarity) between $h$ and $c$ changes in the direction of $V_{i}$.
In the special case that $h$ is the identity function, we obtain an even simpler interpretation of CDAM due to $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$. Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Section A.1 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
@@
We note that introducing $h$ attention heads does not change the above derivations qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section A.1 reads
@@
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section A.1 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
We can also show that CDAM scales the attention by the relevance of the corresponding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT. Analogously to the above single-head discussion we have
@@
A.2 Results for a random subset of ImageNet
Figure S1: Box sensitivity curves obtained from the random ImageNet subset
Figure S2: MIF/LIF perturbation curves obtained from the random ImageNet subset.
[…]
CDAM 801 1423 623 11.3
Integrated CDAM 696 1459 753 12.9
Smooth CDAM 591 1482 890 11.9
Rel. Prop. 475 1432 957 12.7
Partial LRP 654 1329 674 8.4
Att. map 580 1352 771 9.1
SmoothGrad 796 786 -10 0.2
IntGrad 411 1055 644 8.3
Input x Grad 470 935 465 3.1
Table S1: Area under the curve for MIF, LIF and LIF-MIF curves (Fig. S2) and box sensitivity curves (Fig. S1) obtained from the random ImageNet subset.
A.3 Examples of CDAM for a known class on the ImageNet
Figure S3: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S4: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S5: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
A.4 Examples for CDAM on the Lung Image Database Consortium (LIDC)
Figure S6: Examples for the Attention Map and CDAM based on a ViT malignancy model trained on the lung nodule data from LIDC.
Figure S7: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.
Figure S8: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.